RC meeting and Run Plan, Thursday December 8th
From SBS wiki
Return to the RC Meeting Calendar
- 1 RC Meeting Information
- 2 Shift Schedule
- 3 Accelerator Status
- 4 Current Status
- 5 Issues Resolved
- 6 Issues remaining
- 7 Short Term Run Plan
- 8 Planned Accesses to the Hall
- 9 Special Requests
- 10 Notices in Effect for Shift Crew
- 11 Data Collection Status
RC Meeting Information
|Current RC:||Eric King [ericking (@jlab.org or @temple.edu)]|
|Time:||Daily at 3:30PM|
|Location:||Counting House Conference Room|
Remote Connection Details
- Join ZoomGov Meeting
- Meeting ID: 161 903 2184
- Passcode: 260715
- Dial by your location
- +1 669 254 5252 US (San Jose)
- +1 646 828 7666 US (New York)
- +1 551 285 1373 US
- +1 669 216 1590 US (San Jose)
- 833 568 8864 US Toll-free
- Find your local number: https://jlab-org.zoomgov.com/u/acj3pTicMx
This RC Week
- Next empty shift is a Target Operator DAY shift on Sunday 12/11.
Upcoming RC Week
- There are seven (7) remaining shifts to be filled until winter shutdown; I'm sure Andrew would like them filled asap so he doesn't have to chase down any shift workers.
- Everything appears to be running smoothly at the moment.
- Hall C is down for about a week due to a vacuum break.
We are currently taking data on 3-He
Beam Restoration (Wednesday)
The experiment was back up and running yesterday by 8PM. :)
- Beam restoration went smoothly with a beam width of ~330 x ~320. 
- Bill's carbon hole analysis looks fine.  
Target Polarization (Since Recovery)
- Most recent (12:40PM) target polarization is 40.7% 
- Previous Measurements:
- Accelerator has agreed to support a Møller measurement during Friday DAY.
- During the last Møller measurement the accelerator didn't get started until about 9:30'ish.
- Target Operator still required.
- Shift leader should still come in as of now and plan on up to an hour of data taking.
Beam ramp step size issue
- I wasn't clear on the resolution of this as I was out of email chains so here is a record of the ELOGs
- Regarding the fix 
"The chosen fix was to adjust the BCM records on iocbcm to force them to issue channel access updates every second. The default was to not send updates when the value was unchanged. This has been done for IBC1H04CRCUR2, IBC3H00CRCUR4, and IBCAD00CRCUR6 which are used by current locks. IBC0L02Current was also changed at the request of I&C."
- Attenuator was working fine 
- Regarding the fix 
GEM Track Efficiency
- Poor GEM track efficiency, alignment needed for analysis. Last spoke with Zeke at around 1:30AM'ish and he will continue working on alignment this morning after collaborating with other GEM experts.    
- Zeke provided a brief summary of the work Sean completed 
- Analysis with the re-alignment 
- BB GEM APV unreachable and cannot be repaired at this time.
Short Term Run Plan
- Continue production run on He3 target, 36 uA
- Use GEnII-noSBSGEMs DAQ config
- Use ps3=0, ps5=6, others =-1
- BBCal threshold at -366 mV, HCal threshold at -19 mV
- NMR measurement every 3 - 4 hours, close to 4 hours is better
Planned Accesses to the Hall
- I presume that repairs for the BB GEM APV will wait until shutdown.
- None of which I'm aware, please email me ericking (@jlab.org or @temple.edu) if there's something to be added.
Notices in Effect for Shift Crew
- Power down SBS, BB, and Detector HV before making an access to the Hall
Data Collection Status
- 74 out of 141 days (~52.5%)
- Total uncorrected accumulated charge on He3: 47.020 C out of 232C (20.26%)
- 2-Pass: 13.467 C
- 3-Pass: 33.553 C (at end of OWL)